The width of a room is 5ft more than 1/4 the length. The perimeter is 230 ft. What is the width of the room??
Answer by checkley77(12569)

W=L/4+5
2L+2W=230
2L+2(L/4+5)=230
2L+L/2+10=230
(2*2L+L)/2=230-10
(4L+L)/2=220
5L=220*2
5L=440
L=440/5
L=88 ANS.
W=88/4+5
W=22+5
W=27 ANS.
PROOF:
2*88+2*27=230
176+54=230
230=230